OpenStudy (anonymous):

What are the first five terms of the pattern with the formula 21n?

OpenStudy (anonymous):

answers first gets medal and fanned and just to let you guys know my other account got hacked

OpenStudy (mathstudent55):

I don't understand the pattern. Can you write the question exactly as you got it?

OpenStudy (anonymous):

that's how I got it I do online schooling and I copy it and past it

OpenStudy (mathstudent55):

Is it \(21^n\) or \(21n\) ?

OpenStudy (anonymous):

2nd one hold @iambatman @Destinymasha @uri @Falco276 @Hero @Yttrium @Lil_Jerome @Okabe86 @uzumakinaruto1

OpenStudy (mathstudent55):

Usually n takes the values of 1, 2, 3, 4, 5, 6, ... Since you are asked for the first five numbers in the pattern, replace n with 1, and evaluate the expression. Then replace n with 2 and evaluate the expression. Then do the same for n = 3, 4, and 5.

OpenStudy (mathstudent55):

21n = 21 * n 1: 21 * 1 = 2: 21 * 2 = 3: 21 * 3 = 4: 21 * 4 = 5: 21 * 5 = Just do those 5 multiplications.